ineedhelp3344:

Are the expressions –0.5(3x + 5) and –1.5x + 2.5 equivalent? Explain why or why not.

AZ:

Do you know the distributive property? a(b + c) = ab + ac

AZ:

Just distribute the -0.5 inside the parenthesis -0.5(3x + 5) = (-0.5*3x) + (-0.5*5) = What do you get when you simplify it?